Unexpected shutdown of vector fridge on 26 09 2012
Magnet temperatures as I found them. Magnets are hot!
Fridge temperatures as I found them. Fridge is hot
IGH front panel as I found it. All valves are shut. Mixture is behind He3 pump and in He4 dump?!
IGH panel again. Notice P3=243! . Most likely mixture.
Pressure gauges as I found them. Probe>1500 mbar and IVC no Ident -> IVC is vented !? Gas in the OVC due to warming of the fridge. Still>1500. There is mixture in the fridge.
IGH front panel after recovery. After recovery pressures don't match previous values + He3 and He4 are mixed in the dumps. Fridge Crash Sept2012 Pressure Logs MG2-MG5.
for details of the fridge dump pressures, but usually they are 990mB & 1153 mB, and we are now getting 985 & 919 mB.
So there is some mixture missing -- maybe this is in the IVC (-- AlexHamilton
- 27 Sep 2012) ?
Switched off both compressors.
Waiting to warm up.
Oleh wrote that he pumped the IVC into the dumps and now the dumps pressures make sense: P6=1160 and P7=989. This is basically the same as the historical values, which says that the extra gas in the IVC may well have been mixture - but how on earth did it get there?
He Leak Detection
Hooked up Leak detector to Aux Port-1, to leak check the still lines and both condensing lines.
*done whilst waiting for fridge to warm up (Magnet temperatures ~ 242K).
*Note: Here we're used the leak detector's rotary pump as a substitute for pump S4.
Result of leak checking:
IVC is holding pressure so no leakage between IVC and fridge lines (still and both condensing lines).
Perhaps it is due to a blockage.
Cleaning Mixture via Fridge Attempt
The following attempts at cycling the mixture was repeated a couple of times, and the same result was obtained every time - inidicating a blockage in the condenser lines.
Pressure Tests (5/8/12)
The following tests were conducted as per Giorgio's suggestion.
Fridge Logs of the Event
Plots of the fridge pressures & temperatures are provided below.
Crash occurred between 8:31am to 8:34am:
- Sudden heat load applied to all Plates of the fridge.
- Flow rate dropped from 177 to 0.
- Mixture was auto dumped in to P6 (He4 dumps rose from 18mbar to >500mbar)
- Mixture also auto dumped into IVC (P3 rose from ~0bar to 250mbar, MG2 and MG4 rose >1500mbar)
Magnet temperatures also increase.
Compressor Logs during that time period show no change - ie power was still on at that time and the compressors were in normal operation:
C:\Liquefier Logs\CompressorState 2012_09_24 1549.txt
C:\Liquefier Logs\CompressorState 2012_09_24 1551.txt
| Timestamp || Low side PSI || High side PSI||In water temp C || Out water temp C||He gas temp C ||Oil temp C ||Motor Amps ||Diode temp K |
| 08:31:18|| 082.7|| 264.5|| 070.5|| 087.8|| 134.6|| 081.7|| 15.0|| NaN|
| 08:36:26|| 090.0|| 273.9|| 070.9|| 088.5|| 135.0|| 082.4|| 15.0|| NaN|
| Timestamp || Low side PSI || High side PSI||In water temp C |
| Out water temp C||He gas temp C |
|Oil temp C |
|Motor Amps |
|Diode temp K |
|08:03:24||083.2|| 268.3||073.0|| 090.0|| 133.0|| 085.3|| 15.0|| NaN|
| 08:08:33||082.7|| 268.4||073.2|| 090.0|| 133.0|| 085.3|| 15.0|| NaN|
Previously the default template was set up so that all pages were visible by the whole world, and we relied upon authors to use access control to limit access to sensitive pages. That hasn't worked, so now pages are by default only visible to QED group members. Delete the following line if this page is meant to be visible by the whole world.
# Set ALLOWTOPICVIEW = QedGroup